Let V be the vector space of all 2×2 matrices over the field R. let W1 = {"\begin{bmatrix}
 x & -x \\
 y & z
\end{bmatrix}"| x,y,z€R} and W2= {"\begin{bmatrix}
 a & b \\
 -a & c
\end{bmatrix}"| a,b,c€R}. What is the dimensions of W1+W2 and W1 intersection W2 as well?


1
Expert's answer
2021-06-17T19:17:18-0400

"W_1 =\{\begin{bmatrix}
 x & -x \\
 y & z
\end{bmatrix}| x,y,z,\in R\}"

Basis "W_1" , for example,

"E_1=\begin{bmatrix}
 1 & -1 \\
 0 & 0
\end{bmatrix},
E_2=\begin{bmatrix}
0 & 0 \\
 1 & 0
\end{bmatrix}, E_3=\begin{bmatrix}
 0 & 0 \\
 0 & 1
\end{bmatrix}"

"dimW_1=3"


"W_2 =\{\begin{bmatrix}
 a & b \\
 -a & c
\end{bmatrix}| a,b,c,\in R\}"

Basis "W_2" , for example,

"E'_1=\begin{bmatrix}
 1 & 0 \\
 -1 & 0
\end{bmatrix},
E'_2=\begin{bmatrix}
0 & 1\\
 0 & 0
\end{bmatrix}, E'_3=\begin{bmatrix}
 0 & 0 \\
 0 & 1
\end{bmatrix}"

"dimW_2=3"


"W_1+W_2:\\
a_1E_1+a_2E_2+a_3E_3+a_4E'_1+a_5E'_2=0\\
\begin{bmatrix}
 a_1+a_4 & -a_1+a_5\\
 a_2-a_4 & a_3
\end{bmatrix}=\begin{bmatrix}
 0 & 0 \\
 0 & 0
\end{bmatrix}"

"a_1+a_4=0\\
-a_1+a_5=0\\
a_2-a_4=0\\
a_3=0"


"a_4=-a_1\\
a_5=a_1\\
a_2=a_4=-a_1\\
a_3=0"

"dim(W_1+W_2)=5\\
dim(W_1\cap W_2)=dimW_1+dimW_2-\\
-dim(W_1+W_2)=3+3-5=1"